OpenStudy (anonymous):

How do you use the properties of lorgarithms to solve for x: In x + In(x-1)=In 2

OpenStudy (anonymous):

ln x+ln (x-1)=ln2 ln (x)(x-1)=ln2 x^2-x=2 x^2-x-2=0 x^2-2x+x-2=0 x(x-2)+1(x-2)=0 (x-2)(x+!)=0 Either x-2=0,x=2 Or x+1=0,x=-1(Rejected ,because x>or=0,for log x to be defined)